# Asymptotia of Kerr–de Sitter Black Holes

### v1

@@ II Background
Conformal fundamental forms are weighted, rank 2 tensors that generalize extrinsic hypersurface curvatures to conformally embedded hypersurfaces [23]. They are related to the asymptotic stress-energy tensor via the Einstein field equations [24]. The first conformal fundamental form is the induced conformal class of metrics $\overline{\bm{g}}$ on $\Sigma$. For 4-dimensional conformal manifolds, the second through fourth fundamental forms are of interest because they capture the initial data for the conformal Einstein equations. They are given by the trace-free extrinsic curvature $\hskip 1.13809pt\mathring{\rm{II}}{}_{ab}:={\rm II}_{ab}-H\overline{g}_{ab}$, where $H$ is the mean curvature,
${\mathring{{\bf\rm III}}{\hskip 0.56905pt}}{}_{ab}:=W_{\hat{n}ab\hat{n}}-\hskip 1.13809pt\mathring{\rm{II}}{}_{a}{}^{c}\hskip 1.13809pt\mathring{\rm{II}}{}_{bc}+\tfrac{1}{3}\overline{g}_{ab}\hskip 1.13809pt\mathring{\rm{II}}{}_{cd}\hskip 1.13809pt\mathring{\rm{II}}{}^{cd},$
@@
${\mathring{{\bf\rm IV}}{\hskip 0.56905pt}}{}_{ab}:=C^{\top}_{(a|\hat{n}|b)}-\overline{\nabla}^{c}W^{\top}_{\hat{n}(ab)c}+HW_{\hat{n}a\hat{n}b}\,,$
where $\hat{n}$ is the timelike unit normal vector to the boundary $\Sigma$, $\overline{\nabla}$ the Levi-Civita connection of $\overline{g}$ and $\top$ denotes projection along $\Sigma$. The fourth conformal fundamental form ${\mathring{{\bf\rm IV}}{\hskip 0.56905pt}}{}$ is of utmost importance in dimension 4. In Euclidean signature, it is the (locally undetermined) image of the Dirichlet-to-Neumann map for the Poincaré-Einstein problem [25, 26]. In Lorentzian signature, ${\mathring{{\bf\rm IV}}{\hskip 0.56905pt}}{}$ is a divergence-free and trace-free tensor that is part of the free data for Friedrich’s vacuum conformal Einstein field equations [27, 28]. One can show that it is equal to the electric part of the rescaled Weyl tensor in that setting. When matter fields are present, the divergence of ${\mathring{{\bf\rm IV}}{\hskip 0.56905pt}}{}$ no longer vanishes, but instead can be expressed in terms of the stress-energy tensor. Similar relations hold for conformal boundaries of regularized spacetimes with initial isotropic singularities [29]. These underlie cosmological models that extend the causal structure beyond the Big Bang, such as conformal cyclic cosmology [30, 31, 32, 33, 34].
@@ III Asymptotic Generalized Symmetry
and $\sigma$ is a defining density for $\Sigma$. A condition $\mathcal{C}_{abc}=0$ is equivalent to $Q_{ab}$ being a closed $2-$form with respect to the metric $\sigma^{-2}\bm{g}_{ab}$.
@@
The connection $\nabla^{\mathcal{T}}$ can be upgraded to the so-called Thomas-D operator, which maps tractors of weight $w$ to tractors of weight $w-1$ and is given by
@@
̊$\begin{split}\mathring{\top}\nabla_{\hat{n}}^{k}\mathcal{Q}_{abc}&\propto\mathring{\underline{\overline{\rm{k+2}}}}_{{\hskip 0.85358pt}a[b}Q_{c]_{\circ}\hat{n}}^{\top}+\cdots,\\ \mathring{\top}\hat{n}^{c}\nabla_{\hat{n}}^{k}\mathcal{Q}_{(ab)c}&\propto\mathring{\underline{\overline{\rm{k+2}}}}_{{\hskip 0.85358pt}(a}{}^{c}Q^{\top}_{b)_{\circ}c}+\cdots\,.\end{split}$ (11)
In the above $\nabla^{k}_{\hat{n}}:=\hat{n}^{a_{1}}...\hat{n}^{a_{k}}\nabla_{a_{1}}...\nabla_{a_{k}}$, $\propto$ denotes equality up to a non-vanishing constant, and $\mathring{\underline{\overline{\rm{k+2}}}}$ denotes the $(k+2)$th conformal fundamental form (as long as it is defined [40]). This can be used to constrain $\mathring{\underline{\overline{\rm{k+2}}}}$ in terms of the projections of the asymptotic CKY 2-form on the conformal boundary. However, the coefficients of those leading terms in (10) depend on the dimension $d$ and order $k$ and vanish in certain cases. In particular, all of them are 0 when $d=4$ and $k=2$. This limits the utility of these operators at and beyond that order [42]. Note that this problem arises precisely when the conformal fundamental form appearing in (11) is ${\mathring{{\bf\rm IV}}{\hskip 0.56905pt}}{}$, the free data for the conformal Einstein field equations in four dimensions.
@@
$\begin{split}&2{\mathring{{\bf\rm III}}{\hskip 0.56905pt}}{}_{a[b}Q^{\top}{}_{c]\hat{n}}+\overline{g}_{a[b}{\mathring{{\bf\rm III}}{\hskip 0.56905pt}}{}_{c]}{}^{d}Q^{\top}_{d\hat{n}}+\ldots\,,\\ \vspace{0.2cm}&Q^{\top}{}_{(a}{}^{c}{\mathring{{\bf\rm III}}{\hskip 0.56905pt}}{}_{b)c}+\ldots\,,\\ \vspace{0.2cm}&\top\nabla_{\hat{n}}^{2}Q_{ab}-2Q^{\top}{}_{[a}{}^{c}{\mathring{{\bf\rm III}}{\hskip 0.56905pt}}{}_{b]c}+\ldots\,,\\ \vspace{0.2cm}&\top\hat{n}^{b}\nabla_{\hat{n}}^{2}Q_{ab}+{\mathring{{\bf\rm III}}{\hskip 0.56905pt}}{}_{a}{}^{b}Q^{\top}{}_{b\hat{n}}+\ldots\,,\end{split}$
where the subleading terms denoted by $\ldots$ depend on ${\rm II}$, projections of $Q$, and curvatures intrinsic to $\Sigma$. The second and third lines above come from the symmetric and antisymmetric parts of the two-index operator from Equations (10).
@@ IV Rotating Black Holes and Generalized Symmetries
$\begin{split}\nabla_{a}\nabla_{b}\tilde{\sigma}&+\tilde{\sigma}P_{ab}-\tfrac{1}{4}\bm{g}_{ab}(\Delta\tilde{\sigma}+J\tilde{\sigma})=\tfrac{\tilde{\sigma}}{2}\mathring{\widetilde{T{\hskip 0.85358pt}}}_{ab}\,,\\ &I^{2}_{\tilde{\sigma}}=-1+\tfrac{\tilde{\sigma}^{2}}{12}\widetilde{T}\,.\end{split}$ (12)
where $\tilde{\sigma}$ is a defining density for the conformal boundary $\Sigma$, $\widetilde{T}:=\widetilde{T}_{a}{}^{a}$, and the singular physical metric $\widetilde{g}=\tilde{\sigma}^{-2}\bm{g}$. We want to focus on solutions to Equations (12) asymptotic to the Kerr–de Sitter metric in a neighborhood of $\Sigma$. To achieve this we assume that the stress-energy tensor $\widetilde{T}_{ab}$ vanishes on $\Sigma$ and that there exists a solution of Equations (8) and (9) – an asymptotic CKY $2-$form. It follows that,
$\nabla_{a}\nabla_{b}\tilde{\sigma}+\tilde{\sigma}P_{ab}-\tfrac{1}{4}\bm{g}_{ab}(\Delta\tilde{\sigma}+J\tilde{\sigma})$ $=\mathcal{O}(\tilde{\sigma}^{\ell}),$ (13)
@@ IV.1 Solving Geometric Constraints
$\hskip 1.13809pt\mathring{\rm{II}}{}\stackrel{{\scriptstyle\Sigma}}{{=}}0\stackrel{{\scriptstyle\Sigma}}{{=}}{\mathring{{\bf\rm III}}{\hskip 0.56905pt}}{}\,,$
where =${\stackrel{{\scriptstyle\Sigma}}{{=}}}$ denotes equality on $\Sigma$. Proceeding order-by-order, we now derive further constraints arising from Equations (15) and (16). The latter implies that the projected part of the asymptotic CKY tensor vanishes along the conformal infinity, i.e.
@@
General manifestly conformally invariant constraint equations could have been derived, had we not fixed $\hskip 1.13809pt\mathring{\rm{II}}{}\stackrel{{\scriptstyle\Sigma}}{{=}}0\stackrel{{\scriptstyle\Sigma}}{{=}}Q^{\top}$.
@@
$(\nabla_{\hat{n}}^{2}Q_{ab})^{\top}\stackrel{{\scriptstyle\Sigma}}{{=}}$ $\ 2Q^{\top}{}_{\hat{n}[a}\overline{\nabla}_{b]}H-H\overline{\nabla}_{[a}Q^{\top}_{|\hat{n}|b]}\,,$
@@
Upon differentiating twice in the normal direction, we arrive at equations relating $(\nabla_{\hat{n}}^{3}Q_{ab})^{\top}|_{\Sigma}$ and $\hat{n}^{b}\overline{g}_{a}{}^{c}\nabla_{\hat{n}}^{3}Q_{bc}|_{\Sigma}$ to the geometry of the conformal boundary and $Q|_{\Sigma}$, as well as an algebraic constraint for ${\mathring{{\bf\rm IV}}{\hskip 0.56905pt}}{}$,
$\ 2{\mathring{{\bf\rm IV}}{\hskip 0.56905pt}}{}_{a[b}Q^{\top}{}_{|\hat{n}|c]}+\overline{g}_{a[b}{\mathring{{\bf\rm IV}}{\hskip 0.56905pt}}{}_{c]}{}^{d}Q_{\hat{n}d}^{\top}\stackrel{{\scriptstyle\Sigma}}{{=}}0\,.$ (18)
It should be noted that in the setting of asymptotically de Sitter spacetimes one can only derive a differential constraint relating a divergence of ${\mathring{{\bf\rm IV}}{\hskip 0.56905pt}}{}$ to the stress-energy tensor, see e.g. [24].
@@
We now focus on the algebraic equation (18). Given that $Q|_{\Sigma}=Q_{KdS}|_{\Sigma}$ (see Equation (4)) and the induced metric $\overline{g}$ on $\Sigma$ from Equation (5) we may solve it for ${\mathring{{\bf\rm IV}}{\hskip 0.56905pt}}{}$, yielding
$\begin{split}{\mathring{{\bf\rm IV}}{\hskip 0.56905pt}}{}=&A\big{[}\mathrm{d}t^{2}-2a\sin^{2}\theta\mathrm{d}t\mathrm{d}\phi-\tfrac{(1+a^{2})^{2}}{2(1+a^{2}\cos^{2}\theta)}\mathrm{d}\theta^{2}\\ &-\tfrac{1}{2}(1+a^{2}-3a^{2}\sin^{2}\theta)\sin^{2}\theta\mathrm{d}\phi^{2}\big{]}\,,\end{split}$ (19)
@@
A spacetime satisfying the constraints above is asymptotically Kerr–de Sitter in a neighborhood of $\Sigma$. This allows us to construct an asymptotic Killing vector field from the divergence of the asymptotic CKY 2-form $Q_{ab}$ [44]. To achieve that, we need the physical metric and the CKY tensor to orders $-1$ and $-2$ in the defining function, respectively. Given a metric representative of the conformal infinity, a Fefferman–Graham expansion [45] provides an explicit form of an asymptotically vacuum Einstein spacetime metric using the Gaussian normal coordinate to the boundary; see e.g. [46] for the derivation of the Fefferman-Graham form of the rotating black hole metric with negative cosmological constant. Hence, given $\overline{g}$ from Equation (5), the physical metric can be written as,
@@
where $\overline{P}$ is the Schouten tensor of the boundary metric $\overline{g}$. Note that (20) can only be used to recover the exact Kerr–de Sitter solution after inserting appropriate Neumann data at cubic order in $s$.
@@
~$\widetilde{\nabla}_{a}\widetilde{Q}^{ab}=3(1+a^{2})^{2}\delta_{t}^{b}+\mathcal{O}(s)\,.$
@@ IV.2 Matter
$\tau_{ab}=-\hat{n}_{a}\hat{n}_{b}\tau+\mathcal{O}(\tilde{\sigma})\,.$
We also have the following relation between $\tau_{ab}|_{\Sigma}$ and ${\mathring{{\bf\rm IV}}{\hskip 0.56905pt}}{}_{ab}$ [24],
¯$\overline{\nabla}^{b}{\mathring{{\bf\rm IV}}{\hskip 0.56905pt}}{}_{ab}\stackrel{{\scriptstyle\Sigma}}{{=}}\lim_{\tilde{\sigma}\to 0}\tilde{\sigma}^{-1}\tau_{\hat{n}a}^{\top}-\tfrac{1}{3}\overline{\nabla}_{a}\tau\,,$ (21)
where the limit on the right-hand side is obviously well defined. The fourth conformal fundamental form ${\mathring{{\bf\rm IV}}{\hskip 0.56905pt}}{}$ is given by (19), so we can now use (21) to constrain the free function $A$, yielding

### v2

# Asymptopia of Kerr–de Sitter Black Holes

@@ II Background
Conformal fundamental forms are weighted, rank 2 tensors that generalize extrinsic hypersurface curvatures to conformally embedded hypersurfaces [23]. They are related to the asymptotic stress-energy tensor via the Einstein field equations [24]. The first conformal fundamental form is the induced conformal class of metrics $\overline{\bm{g}}$ on $\Sigma$. For 4-dimensional conformal manifolds, the second through fourth fundamental forms are of interest because they capture the initial data for the conformal Einstein equations. They are given by the trace-free extrinsic curvature $\hskip 1.13809pt\mathring{\rm{I\hskip-0.56905ptI}}{}_{ab}:={\rm I\hskip-0.56905ptI}_{ab}-H\overline{g}_{ab}$, where $H$ is the mean curvature,
${\mathring{{\bf\rm I\hskip-0.56905ptI\hskip-0.56905ptI}}{\hskip 0.56905pt}}{}_{ab}:=W_{\hat{n}ab\hat{n}}-\hskip 1.13809pt\mathring{\rm{I\hskip-0.56905ptI}}{}_{a}{}^{c}\hskip 1.13809pt\mathring{\rm{I\hskip-0.56905ptI}}{}_{bc}+\tfrac{1}{3}\overline{g}_{ab}\hskip 1.13809pt\mathring{\rm{I\hskip-0.56905ptI}}{}_{cd}\hskip 1.13809pt\mathring{\rm{I\hskip-0.56905ptI}}{}^{cd}$,
@@
${\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}}{}_{ab}:=C^{\top}_{(a|\hat{n}|b)}-\overline{\nabla}^{c}W^{\top}_{\hat{n}(ab)c}+HW_{\hat{n}a\hat{n}b}\,$,
where $\hat{n}$ is the timelike unit normal vector to the boundary $\Sigma$, $\overline{\nabla}$ the Levi-Civita connection of $\overline{g}$ and $\top$ denotes projection along $\Sigma$. The fourth conformal fundamental form $\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}$}{} is of utmost importance in dimension 4. In Euclidean signature, it is the (locally undetermined) image of the Dirichlet-to-Neumann map for the Poincaré-Einstein problem [25, 26]. In Lorentzian signature, $\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}$}{} is a divergence-free and trace-free tensor that is part of the free data for Friedrich’s vacuum conformal Einstein field equations [27, 28]. One can show that it is equal to the electric part of the rescaled Weyl tensor in that setting. When matter fields are present, the divergence of $\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}$}{} no longer vanishes, but instead can be expressed in terms of the stress-energy tensor. Similar relations hold for conformal boundaries of regularized spacetimes with initial isotropic singularities [29]. These underlie cosmological models that extend the causal structure beyond the Big Bang, such as conformal cyclic cosmology [30, 31, 32, 33, 34].
@@ III Asymptotic Generalized Symmetry
and $\sigma$ is a defining density for $\Sigma$. A condition $\mathcal{C}_{abc}=0$ is equivalent to $Q_{ab}$ being a closed 22-form with respect to the metric $\sigma^{-2}\bm{g}_{ab}$.
@@
The connection $\nabla^{\mathcal{T}}$ can be upgraded to the so-called Thomas-D operator, which maps tractors of weight $w$ to tractors of weight w1w-1 and is given by
@@
$\begin{split}\mathring{\top}\nabla_{\hat{n}}^{k}\mathcal{Q}_{abc}&\propto\mathring{\underline{\overline{\rm{k+2}}}}_{{\hskip 0.85358pt}a[b}Q_{c]_{\circ}\hat{n}}^{\top}+\cdots,\\ \mathring{\top}\hat{n}^{c}\nabla_{\hat{n}}^{k}\mathcal{Q}_{(ab)c}&\propto\mathring{\underline{\overline{\rm{k+2}}}}_{{\hskip 0.85358pt}(a}{}^{c}Q^{\top}_{b)_{\circ}c}+\cdots\,.\end{split}$ (11)
In the above $\nabla^{k}_{\hat{n}}:=\hat{n}^{a_{1}}...\hat{n}^{a_{k}}\nabla_{a_{1}}...\nabla_{a_{k}}$, $\propto$ denotes equality up to a non-vanishing constant, and $\mathring{\underline{\overline{\rm{k+2}}}}$ denotes the $(k+2)$th conformal fundamental form (as long as it is defined [40]). This can be used to constrain $\mathring{\underline{\overline{\rm{k+2}}}}$ in terms of the projections of the asymptotic CKY 2-form on the conformal boundary. However, the coefficients of those leading terms in (10) depend on the dimension $d$ and order $k$ and vanish in certain cases. In particular, all of them are 0 when $d=4$ and $k=2$. This limits the utility of these operators at and beyond that order [42]. Note that this problem arises precisely when the conformal fundamental form appearing in (11) is $\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}$}{}, the free data for the conformal Einstein field equations in four dimensions.
@@
[…]
where the subleading terms denoted by $\ldots$ depend on $\rm I\hskip-0.56905ptI$}, projections of $Q$, and curvatures intrinsic to $\Sigma$. The second and third lines above come from the symmetric and antisymmetric parts of the two-index operator from Equations (10).
@@ IV Rotating Black Holes and Generalized Symmetries
[…]
where $\tilde{\sigma}$ is a defining density for the conformal boundary $\Sigma$, $\widetilde{T}:=\widetilde{T}_{a}{}^{a}$, and the singular physical metric $\widetilde{g}=\tilde{\sigma}^{-2}\bm{g}$. We want to focus on solutions to Equations (12) asymptotic to the Kerr–de Sitter metric in a neighborhood of $\Sigma$. To achieve this we assume that the stress-energy tensor $\widetilde{T}_{ab}$ vanishes on $\Sigma$ and that there exists a solution of Equations (8) and (9) – an asymptotic CKY 22-form. It follows that,
[…]
@@ IV.1 Solving Geometric Constraints
$\hskip 1.13809pt\mathring{\rm{I\hskip-0.56905ptI}}{\hskip 0.0pt}\stackrel{{\scriptstyle\Sigma}}{{=}}0\stackrel{{\scriptstyle\Sigma}}{{=}}{\mathring{{\bf\rm I\hskip-0.56905ptI\hskip-0.56905ptI}}{\hskip 0.56905pt}}{}\,$,
where $\stackrel{{\scriptstyle\Sigma}}{{=}}$} denotes equality on $\Sigma$. Proceeding order-by-order, we now derive further constraints arising from Equations (15) and (16). The latter implies that the projected part of the asymptotic CKY tensor vanishes along the conformal infinity, i.e.
@@
General manifestly conformally invariant constraint equations could have been derived, had we not fixed $\hskip 1.13809pt\mathring{\rm{I\hskip-0.56905ptI}}{\hskip 0.0pt}\stackrel{{\scriptstyle\Sigma}}{{=}}0\stackrel{{\scriptstyle\Sigma}}{{=}}Q^{\top}$.
@@
$(\nabla_{\hat{n}}^{2}Q_{ab})^{\top}\stackrel{{\scriptstyle\Sigma}}{{=}}$ $\ 2Q^{\top}{}_{\hskip-5.69046pt\hat{n}[a}\overline{\nabla}_{b]}H-H\overline{\nabla}_{[a}Q^{\top}_{|\hat{n}|b]}\,$,
@@
Upon differentiating twice in the normal direction, we arrive at equations relating $(\nabla_{\hat{n}}^{3}Q_{ab})^{\top}|_{\Sigma}$ and $\hat{n}^{b}\overline{g}_{a}{}^{c}\nabla_{\hat{n}}^{3}Q_{bc}|_{\Sigma}$ to the geometry of the conformal boundary and $Q|_{\Sigma}$, as well as an algebraic constraint for $\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}$}{},
$\ 2{\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}}{}_{a[b}Q^{\top}{}_{\hskip-5.69046pt|\hat{n}|c]}+\overline{g}_{a[b}{\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}}{}_{c]}{}^{d}Q_{\hat{n}d}^{\top}\stackrel{{\scriptstyle\Sigma}}{{=}}0\,.$ (18)
It should be noted that in the setting of asymptotically de Sitter spacetimes one can only derive a differential constraint relating a divergence of $\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}$}{} to the stress-energy tensor, see e.g. [24].
@@
We now focus on the algebraic equation (18). Given that $Q|_{\Sigma}=Q_{KdS}|_{\Sigma}$ (see Equation (4)) and the induced metric $\overline{g}$ on $\Sigma$ from Equation (5) we may solve it for $\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}$}{}, yielding
[…]
@@
A spacetime satisfying the constraints above is asymptotically Kerr–de Sitter in a neighborhood of $\Sigma$. This allows us to construct an asymptotic Killing vector field from the divergence of the asymptotic CKY 2-form $Q_{ab}$ [44]. To achieve that, we need the physical metric and the CKY tensor to orders 1-1 and 2-2 in the defining function, respectively. Given a metric representative of the conformal infinity, a Fefferman–Graham expansion [45] provides an explicit form of an asymptotically vacuum Einstein spacetime metric using the Gaussian normal coordinate to the boundary; see e.g. [46] for the derivation of the Fefferman-Graham form of the rotating black hole metric with negative cosmological constant. Hence, given $\overline{g}$ from Equation (5), the physical metric can be written as,
@@
where $\overline{P}$ is the Schouten tensor of the boundary metric $\overline{g}$. Note that (20) can only be used to recover the exact Kerr–de Sitter solution after inserting appropriate Neumann data at cubic order in $s$ (see e.g. [47] for the form of that data). All higher-order terms in the Fefferman–Graham expansion are then determined by the boundary metric and the cubic order term. In the next section, we will constrain the trace-free part of the Neumann data for asymptotically Kerr–de Sitter spacetimes.
@@
$\widetilde{\nabla}_{a}\widetilde{Q}^{ab}=3(1+a^{2})^{2}\delta_{t}^{b}+\mathcal{O}(s)\,.$ (21)
@@ IV.2 Matter
[…]
We also have the following relation between $\tau_{ab}|_{\Sigma}$ and ${\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}}{}_{ab}$ [24],
$\overline{\nabla}^{b}{\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}}{}_{ab}\stackrel{{\scriptstyle\Sigma}}{{=}}p_{a}-\tfrac{1}{3}\overline{\nabla}_{a}\tau\,$, (22)
where $p_{a}:=\lim_{\tilde{\sigma}\to 0}\tilde{\sigma}^{-1}\tau_{\hat{n}a}^{\top}$ is obviously well-defined. The fourth conformal fundamental form $\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}$}{} is given by (19), so we can now use (22) to constrain the free function $A$, yielding
@@
IV.3 Gravitational Charges
We will now provide a partial interpretation the of function $A$ in the admissible free data $\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}$}{} from (19) for the asymptotically Kerr–de Sitter spacetimes. This will be done by expressing quasi-local charges on the conformal boundary introduced in [21] in terms of that function (see [47] for comparison with alternative definitions of charges in this setting). A gravitational charge $C_{\xi}[S]$ associated with a conformal Killing vector $\xi^{a}$ and a surface $S$ embedded in the conformal boundary is defined as
$C_{\xi}[N]:=\frac{1}{8\pi}\oint\displaylimits_{N}\left({\mathring{{\bf\rm I\hskip-0.56905ptV}}{\hskip 0.56905pt}}{}_{ab}+\tfrac{1}{3}\tau\overline{g}_{ab}\right)\xi^{a}\hat{r}^{b}\mathrm{d}N\,$,
where $\hat{r}^{a}$ and $dN$ are the unit normal to $N$ in $\Sigma$ and the area element of that surface with respect to metric representative $\overline{g}$, respectively. A simple calculation shows that the quantity above is conformally invariant.
The charges corresponding to (conformal) Killing vectors $\partial_{t}$ and $\partial_{\phi}$ on the conformal boundary of the exact Kerr–de Sitter metric are independent of the surface $S$ and are proportional to mass and angular momentum parameters. In the current setting, those two vectors are also infinitesimal generators of symmetries of the conformal boundary (where the first one arises from the divergence of the asymptotic CKY tensor via Equation (21)). Hence, we can treat the corresponding charges as definitions of quasi-local mass and angular momentum of asymptotically Kerr–de Sitter spacetime. In particular, for the $t=\mathrm{const}$ spheres $S_{t}$ in the metric representative $\overline{g}$ from (5) we arrive at
$C_{\partial_{t}}[S_{t}]=\frac{1}{8\pi}\oint\displaylimits_{S^{2}}\left(A+\frac{\tau}{3(1+a^{2})^{2}}\right)\,\mathrm{d}S^{2}$
and
$C_{\partial_{\phi}}[S_{t}]=-\frac{3a}{16\pi}\oint\displaylimits_{S^{2}}A\sin^{2}\theta\,\mathrm{d}S^{2}$
where the integrals on the right-hand sides are done over the unit 22-sphere equipped with the standard round metric, i.e. $\mathrm{d}S^{2}=\sin\theta\mathrm{d}\theta\mathrm{d}\phi$.
In the case of $A=\frac{2m}{(1+a^{2})^{2}}$ and $\tau=0$ we recover Kerr-de Sitter charges of [21]. Our expressions can then be thought as generalizations of (quasi-local) mass and angular momentum to the current scenario, with the relation between the function $A$ determining free data and those physical quantities.